OpenStudy (anonymous):

What is the similarity ratio of a cube with volume 2197 m3 to a cube with volume 512 m3?

jimthompson5910 (jim_thompson5910):

if the first cube has a volume of 2197 cubic meters, then what is the length of the side of the cube?

OpenStudy (anonymous):

Uhh.. 2197?

OpenStudy (anonymous):

\[\sqrt[3]{2197} = 13\]

OpenStudy (anonymous):

\[\sqrt[3]{512} = 8\], what is the ratio of 13 to 8? :)

jimthompson5910 (jim_thompson5910):

solve for s V = s^3 2197 = s^3 .... .... .... s = ???

OpenStudy (anonymous):

13:8?

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

Thanks, both of you (: